OpenStudy (anonymous):

y=a(x-h)^2+k what do the letters represent and do?

hero (hero):

(h,k) is the vertex (x,y) is any point on the parabola a is the vertical stretch/compression factor

OpenStudy (anonymous):

does h determine if it is a positive or negative parabola?

hero (hero):

If a is positive, then the parabola opens up. If a is negative, then the parabola opens down.

OpenStudy (anonymous):

i'm sorry but what is the job of h and k then ? i know itz the vertex just a little confused ..

hero (hero):

(h , k) is just the vertex. The vertex is just the peak or valley of the graph of a parabola. The equation is written in vertex form to make it convenient to identify where the vertex is on the parabola. I don't really know what more you need about it.

OpenStudy (anonymous):

mmm... ok thank-you :) i'll try my best from the information given :D

hero (hero):

Well, maybe this is what you need: The parent graph is \(y = x^2\). \(h\) shifts \(x^2\) horizontally h units to the right \(k\) shifts \(x^2\) vertically h units upwards. You end up with \(y = (x - h)^2 + k\) And of course (h,k) together represent the vertex.

hero (hero):

Well the above assumes that \(a = 1\) . If \(a\) is something other than 1, then the graph will be different of course.

hero (hero):

You have more than enough information to understand this I believe.

OpenStudy (anonymous):

perfect! now i understand thank-you so much!

hero (hero):

Well, I'm going to re-write the above. If you don't have \(h\) or \(k\) , then you just have \(y = ax^2\) And then \(h\) shifts the graph horizontally left or right depending on the value of h \(k\) shifts the graph vertically up or down depending on its value. After the shifting is finished, you have \(y = a(x - h)^2 + k\)

OpenStudy (anonymous):

got it thanks! (:

hero (hero):

When I said "you don't have \(h\) or \(k\) I meant that \((h,k) = (0,0)\). Every parabola must have an \(h\) and \(k\) value. Even if their values are both zero.

OpenStudy (anonymous):

ohh ok! so when grahing it would strat at (0,0) and in the equation i dont really have to include it correct?

hero (hero):

When you are graphing a parabola you should start with the vertex if you know what the vertex is. If you have \(y = 5x^2\) obviously the vertex is (0,0) If you have \(y = 5(x - 4)^2 + 7\) then the vertex is (4,7) and you would start with (4,7) not (0,0)

OpenStudy (anonymous):

it's fine if u dont know but imma ask maybe u will understand. if the parabola is not a function but it is a "sleeping" porabula will it be the same equation?

hero (hero):

I know what the question is asking. The graph will not be the same.

OpenStudy (anonymous):

so itz a different equation since itz no longer the same?

hero (hero):

It has to be different since its graph has changed. Every graph has a unique equation associated with it.

OpenStudy (anonymous):

alrighty thanks